5. Fur colour in mice is a single gene trait controled by two alleles. In a population of 75 mice, 21 are homozygous dominant, 37 are heterozygous dominant, and 17 are homozygous recessive. What is the frequency of the dominant allele in the population? Show all work and record your answer as a value between O and 1 rounded to two decimal places 6. The Hardy-Weinberg principle states that allele and genotype frequencies remain constant from one generation to the next, as long as specific conditions are met. 2 Write Yes or No for the conditions that must be met from the provided statements. Mutations are exponentially occurring. All members of the population breed. Everyone produces the same number of offspring. The population is infinitely large. There is no migration in or out of the population. Nonet mutations are occurring. Natural selection of beneficial traits is occurring. Natural selection is not occurring. All mating is completely random. Offspring are able to migrate out of the population
3. In garden pea plants, the yellow seed colour (n is dominant over the In a population of 24 pea plants, 15 pea plants are homozygous d heterozygous dominant, and 3 are recessive. seed colour . dominant, 6 are Show all work and record your response Find the frequency of the following. significant digits. using three b. Yy genotype:_ yy genotype: C. In fruit flies, long wings (L) are dominant over short wings 4. の. In a population of 45 fruit flies, 27 have LL genotypes, 13 have LI genotypes, and 5 havelI genotypes. Find the frequency of the following. Show all work and record your answer as a value between 0 and 1 rounded to two decimal places. long-winged fruit flies: a. b. short-winged fruit flies:
Match the following terms with their correct definitions listed below 1. Definitions Terms Gene pool A group of individuals from the same species living in the same region at the same time A. Genes A population that maintains its allele frequencies by meeting the five given conditions B. Alleles Hardy-Weinberg Equilibrium C. The genetic composition of a given population Microevolution Different forms of the same gene D. Population E. A gradual change in the allele frequencies F DNA that code for specific proteins Choose the correct frequency for the following statements. 2. Frequency Statement Asample of a mouse population has a frequency of 0.18 for the dominant allele for black hair coat (B). A. Allele frequency B. Genotype frequency C. Phenotype frequency Asample of a mouse population has a frequency of 0.32 for heterozygote coat colour individuals (Bb). A sample of a mouse population has a frequency of 0.48 for white coat colour individuals.
Answer #1

5.

Total no. Of progeny = 75

No. Of homozygous dominant = 21

No. Of heterozygous dominant = 37

No. Of homozygous recessive = 17

Genotypic frequency of homozygous dominant = 21/75 = 0.28

Allelic frequency of dominant allele = √0.28 = 0.529

Genotypic frequency of homozygous recessive = 17/75 = 0.226

Allelic frequency of recessive allele = 0.475

6.

As per HWE, a population should be non evolving, no evolutionary force must act on the population.

◆ no mutation

◆ no migration

◆ random mating

◆ large population

◆ no selection

◆ large breeding population

3. Total pea plants = 24

Homozygous dominant = 15

Heterozygous dominant = 6

Homozygous recessive = 3

Genotypic frequency of homozygous dominant = 15/24 = 0.625

Allelic frequency of dominant allele = √0.625 = 0.791

Genotypic frequency of homozygous recessive = 3/24 = 0.125

Allelic frequency of recessive allele = √0.125 = 0.354

Genotypic frequency of heterozygous dominant = 6/24 = 0.25

4. Total fruit flies = 45

Long winged fruit flies = LL + Ll = 27 + 13 = 40

Genotypic frequency of long winged fruit flies = 40/45 = 0.889

Short winged fruit flies = ll = 5

Genotypic frequency of short winged fruit flies = 5/45 = 0.112

If you find the answer useful then please hit the like button.